OpenStudy (anonymous):

A triangle with sides lengths 7, 9 and 11 is a right triangle. True or False

OpenStudy (aravindg):

What did you get ?

OpenStudy (anonymous):

yes I believe it is False

OpenStudy (aravindg):

why false?

OpenStudy (anonymous):

7^2 + 9^2 does not equal 11^2

OpenStudy (anonymous):

\[a^2 + b^2 = c^2 \]

OpenStudy (aravindg):

Good :)

OpenStudy (aravindg):

\[\checkmark\]

OpenStudy (anonymous):

\[7^2 + 9^2 = 11^2 \]

OpenStudy (anonymous):

cool thanks!

OpenStudy (anonymous):

that would be pythagorean theorem which only works with right trianlges...if the 2 smaller sides squared do not equal the larger squared then it can't be a right triangle

OpenStudy (aravindg):

@some_someone really ?

OpenStudy (anonymous):

@some_someone please do not just answer questions. the poitn of open study is to provide learning ..not provide answers

OpenStudy (anonymous):

I did not?

OpenStudy (aravindg):

@Spectrum By the way someone just pulled a fast one :)

OpenStudy (anonymous):

rofl ikr

OpenStudy (aravindg):

lol

OpenStudy (anonymous):

\[49 + 81 = 121\] \[130 = 121 \] FALSE

OpenStudy (anonymous):

so @chocodropa7 you were correct :)

OpenStudy (anonymous):

do you want a treat now someone?

OpenStudy (anonymous):

nope. DO YOU?

OpenStudy (aravindg):

"KEEP CALM AND CALL BATMAN " :)

OpenStudy (anonymous):

lol xD Aravind but your BATMAN, aren't you...? lol